OpenStudy (anonymous):

1. Write using exponents. (–4)(–4)

OpenStudy (amistre64):

multiplication is a more concise way of writing addition, k added n times:\[k+k+k+...+k=nk\] exponents are a more concise way of writing multiplication k multiplied n times... \[k*k*k*k*...*k=k^n\]

OpenStudy (amistre64):

so, first determine you value of k then determine how many times it is being multiplied to itself, n

OpenStudy (anonymous):

O.o

OpenStudy (amistre64):

either you do not understand my explanation, or you perfer others to do your thinking for you ...

OpenStudy (anonymous):

no I don't understand you. and I like to do my own thinking thank you :)

OpenStudy (nincompoop):

look at the patterns: (5)(5)(5) = 5^3 here, your exponent is 3. meaning that 5 has been multiplied three times. (-10)(-10)(-10)(-10)(-10) = (-10)^5 here, your exponent is 5. meaning that (-10) has been multiplied five times. (a)(a) = (a)^2 here, your exponent is 2. meaning that (a) has been multiplied twice. @amistre64 showed you something with k instead of a. these letters are just place holders, meaning they could be any value. it is the same concept, the k has been multiplied by how many times it appeared in the equation. he used n as a placeholder for the exponents to indicate that n will be how many times k will be multiplied.

OpenStudy (anonymous):

ooooo ok :)

OpenStudy (nincompoop):

okay as in you understand now?

OpenStudy (anonymous):

yes thank you!!

OpenStudy (nincompoop):

awesomeness! thank amistre64 :)

OpenStudy (anonymous):

soooo would my answer be 16?

OpenStudy (nincompoop):

correct! negative multiplied by a negative results a positive

OpenStudy (anonymous):

yay!!!!! ^.^ so when I write it down it would be.......